the ten's digit of a two digit number is 3 greater than the unit's digit. if the number is divided by the sum of the digits, the quotient is 6 and the remainder is 8. find the number.
Answer by stanbon(75887)

the ten's digit of a two digit number is 3 greater than the unit's digit. if the number is divided by the sum of the digits, the quotient is 6 and the remainder is 8. find the number.
----
Let the number be 10t+n
----
Equations:
t = u + 3.
(10t+u)/(t+u) = 6 + 8/(t+u)
------------------------------------
Substitute and solve for "u":
(10u + 30)/(u+3+u) = 6 + 8/(u+3+u)
---
(10u + 30)/(2u+3) = 6 + 8/(2u+3)
----
Multiply thru by (2u+3) to get:
10u+30 = 6(2u+3) + 8
10u + 30 = 12u + 26
2u = 4
u = 2
------
Since t = u+3, t = 5
---
The number is 10t+u = 52
================================
cheers,
Stan H.